OpenStudy (anonymous):

The sum of the digits of a two-digit number is 12.If the digits are reversed,the new number is 18 more than the original number.Find the original number.

OpenStudy (anonymous):

x+y = 12 10y+x = 10X + y +18

OpenStudy (anonymous):

where x and y are the 2 digits

OpenStudy (anonymous):

x+y=12 x+10y=y+10x+18 x-y=2. gives u x=7,y=5 the no is 57

OpenStudy (anonymous):

jab ek banda reply kar chuka to kyun sab zor lagate ho yaar?

OpenStudy (anonymous):

TLet the original number be 10x + (12-x) Then on reversing it becomes 10(12-x) + x Atq 10(12-x) + x = 10x + (12-x) + 18 Solve this, get value of x and then get the two numbers